40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 918     Entry time: Tue Aug 10 18:23:46 2010
Author: Jenna 
Type: Laser 
Category: GYRO 
Subject: Error signal investigations 

The cavity still isn't locking properly, so we're sweeping the cavity to have a look at the error signal. The first picture is with everything setup the same way as when we walked in today, and for the second we adjusted the length of a BNC cable and messed with the phase knob on the PDH box to get rid of the DC offset on the error signal.

Blue is the PDH error signal, Yellow is the reflection PD signal, and pink is the waveform we're using to sweep the cavity.

Attachment 1: TEK00001.PNG  15 kB  | Hide | Hide all
TEK00001.PNG
Attachment 2: TEK00003.PNG  15 kB  Uploaded Tue Aug 10 19:26:48 2010  | Hide | Hide all
TEK00003.PNG
ELOG V3.1.3-